ToxicaDennisIce:

Swati is substituting t = 5 and t = 9 to determine if the two expressions are equivalent. 8 (4 t minus 3) 32 t minus 24 Which statement is true?

ToxicaDennisIce:

YNWMelly:

8(4t-3) plug in 5. 8(20-3) = 8(17) = 136. plug in 9. 8(36-3) = 8(33) = 264 this is the first equation

ToxicaDennisIce:

ok

YNWMelly:

32t-24 plug in 5. 32(5) -24 = 160-24 = 136. plug in 9. 32(9) -24 = 288-24 = 264

YNWMelly:

do you get it now?

ToxicaDennisIce:

Both expressions are equivalent to 264 when t = 9.

YNWMelly:

PERFECT!

ToxicaDennisIce:

yayyyyyyyyyy

YNWMelly:

Good Job

ToxicaDennisIce:

medal for u

YNWMelly:

Thanks

ToxicaDennisIce:

ur welcome